OpenStudy (anonymous):

Players X and Y roll a pair of dice alternately. The person who gets 11 first wins. If X starts, what is his chance of winning?

OpenStudy (foolaroundmath):

First find the probability of getting 11. Probability that X wins = Probability that X gets 11 first time + (prob that X doesnt get 11,Y doesnt get 11,X gets 11) + (prob that X gets 11 in his third trial) + ... Its an infinite series which converges.

OpenStudy (shubhamsrg):

we may get 11 by = (5,6) or (6,5) p(11) = 2/36 p(not 11)= 34/36 p(x wins) = p(x gets 11) + p(x doesnt get 11)*p(y doesnt get 11)*p(x gets 11) + p(x doesnt get 11)*p(y doesnt get 11)*p(x doesnt get 11)*p(y doesnt get 11)*p(x gets 11) .... and so on =>2/36 + (34/36)*(34/36)*(2/36) + ........ =>(2/36)(1 + 34/36 + (34/36)^2 ........... ) hope you can simplify further,, and hope my soln was clear..

OpenStudy (anonymous):

Answer is 18/35..not wat the series will give

OpenStudy (anonymous):

how to solve it but??

OpenStudy (shubhamsrg):

sorry a correction there..not 1 + 34/36 + (34/36)^2 ........... but is 1 + (34/36)^2 + (34/36)^4 ........... and its an infinite GP whose common difference is a fraction sum=a/(1-r) a= first term r= common difference still doesnt give 18/35 though..hmm..

OpenStudy (shubhamsrg):

ohh wait,,it does give 18/35..

OpenStudy (shubhamsrg):

:D

OpenStudy (anonymous):

how come... iam not getting it..

OpenStudy (shubhamsrg):

you evaluated the sum ?

OpenStudy (anonymous):

its giving 324/35

OpenStudy (anonymous):

not possible

OpenStudy (shubhamsrg):

note that the whole sum is multiplied by (2/36)

OpenStudy (anonymous):

hey i got it...

OpenStudy (anonymous):

yes i forgot to multiply 2/36.. thanks @shubhamsrg

OpenStudy (shubhamsrg):

nevermind,glad to help :)